THE PHOTOIONIZATION SPECTROSCOPY OF THE ZINC MONOETHYL RADICAL AND ITS POSI-TIVE ION

\underline{MICHAEL B. PUSHKARSKY}, VADIM L. STAKHURSKY, AND TERRY A. MILLER, The Ohio State University, Dept. of Chemistry, Laser Spectroscopy Facility, 120 W. 18th Avenue, Columbus, Ohio 43210.

The high resolution spectroscopy of many closed-shell ions has for years been relatively inaccessible. However, ZEKE spectroscopy applied to open-shell radicals produced in situ in free jets enables one to probe the ground electronic state of the corresponding closed-shell ions. In the recent past we have investigated^{a}^{,}^{b} the ^{1}A_{1} ground electronic state of the cations of three Group IIa and IIb organometallic radicals, MCH_{3}, (M=Mg, Zn, Cd). In this talk we will present the spectroscopic study of the zinc monoethyl radical, and its positive ion. ZnC_{2}H_{5} has been produced in situ in a free jet by laser photolysis of Zn(C_{2}H_{5})_{2}. The vibrational structure of the Ã electronic state of the zinc monoethyl radical and the X̃ state of its cation have been probed by (1+1′) REMPI and ZEKE, respectively. The REMPI spectrum is nearly identical with the previously reported^{c} LIF spectrum. The zinc monoethyl ionization potential has been found to be 56380(5) cm^{−1}. The ZEKE study has revealed complex vibrational structure in the ground electronic state of the ion, dominated by progressions involving ν_{9}, (ω_{9} = 210(3) cm^{−1}), the low frequency in-plane Zn-C-C bending mode. The details of the cation’s vibrational assignment will be given. Implications of our observations for the carbon-metal bonding in the metal alkyl radical molecular systems will be discussed.

aS. E. Panov, D. E. Powers, and T. A. Miller, J. Phys. Chem., 108, 1335 (1998).

bT. A. Barckholtz,D.E. Powers, T.A. Miller, and B. E. Bursten, J. Am. Chem Soc., 121, XXXX (April, 1999).

cI. M. Povey, A. J. Bezant, G. K. Corlett, and A. M. Ellis,J. Phys. Chem., 98, 10427 (1994).